OpenStudy (superhelp101):

Determine the number and type of solutions for x2 + 9x + 7 = 0. 1 Rational Double 2 Irrational 2 Rational 2 Complex

OpenStudy (mathmate):

Can you find the discriminant? For a quadratic equation \(ax^2+by+c\), the discriminant is \(b^2-4ac\)

OpenStudy (superhelp101):

yes the discriminant is 53

OpenStudy (mathmate):

What would be the interpretation if the discriminant is >0 and real, retriceing a,b, & c are real?

OpenStudy (mathmate):

* assuming

OpenStudy (superhelp101):

2 real roots?

OpenStudy (mathmate):

exactly! Good work! :)

OpenStudy (superhelp101):

I'm not sure about rational and irrational thou?

OpenStudy (aum):

If the discriminant is a perfect square you will get two rational roots. If the discriminant is not a perfect square you will get two irrational roots.

OpenStudy (superhelp101):

so 53 isn't so it's option B

OpenStudy (mathmate):

A rational number is a number that can be represented by a fraction of integers. An irrational number is a number that result from square root of non-perfect squares.

OpenStudy (mathmate):

Since the solution involves \(\sqrt{53}\), the solution will be irrational. On the other hand, as @aum said, if the discriminant is a perfect square, then the solution is rational.

OpenStudy (superhelp101):

thank you to both of you!!! :) This one is C I believe If an equation has a positive discriminant, how many times will the graph touch the x-axis? zero 1 2 3

OpenStudy (mathmate):

For \(ax^2+bx+c=0\) the solution is \(\huge \frac{-b\pm \sqrt{b^2-4ac}}{2a}\) so if the discriminant is irrational, the solution is irrational as well.

OpenStudy (superhelp101):

oh I see

OpenStudy (mathmate):

C is correct, can you explain why?

OpenStudy (superhelp101):

I used an example of one of my previous problems. That problem too had two solutions

OpenStudy (mathmate):

Correction to previous post: if the discriminant is not a perfect square, the solution will be irrational. I suppose you caught that! :)

OpenStudy (superhelp101):

:)

OpenStudy (mathmate):

Here the discriminant is positive, so the solution is \(\huge \frac{-b\pm something}{2a}\) so that makes two distinct roots, hence crosses the x-axis twice.

OpenStudy (superhelp101):

yay so I was right! :)

OpenStudy (superhelp101):

I will post another in another post.